# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v2

@@ 6 Discussion
6.1 Reference Model Anchoring: Key to ADPO’s Success
Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$ for implicit KL regularization. The crucial difference lies in how they anchor:
  • Standard DPO: Anchors the difference: $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$
  • ADPO: Anchors each score individually: $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}$ and $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$ separately before comparing
This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—anchoring is trust region by design. Our experiments (Figure 1) demonstrate that ADPO shows relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor for robustness under noise.
6.2 Why Does Hard Outperform Soft Under Heavy Noise (in our controlled setting)?
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios). This appears counter-intuitive. We provide a toy illustration and gradient analysis specific to our noise generation process:
Toy illustration.
Consider a binary Bradley-Terry teacher with observed score difference $\tilde{\Delta R}=\Delta R+\epsilon$, where with probability $p$ an outlier sets $\epsilon\to\pm\infty$ so that $q=\sigma(\beta_{r}\tilde{\Delta R})\approx 0.5$ in expectation across mixed-sign outliers. For Soft-DPO the expected per-pair gradient is
$g_{\text{soft}}=\beta\{\sigma(\beta\Delta_{\theta})-\mathbb{E}[q]\}\approx\beta\{\sigma(\beta\Delta_{\theta})-0.5\},$
which vanishes near $\Delta_{\theta}=0$. In contrast, Hard-DPO draws a Bernoulli label $y\in\{0,1\}$ even when $q\approx 0.5$, yielding
$g_{\text{hard}}=\beta\{\sigma(\beta\Delta_{\theta})-y\},\quad\mathrm{Var}[g_{\text{hard}}]=\beta^{2}\sigma(\beta\Delta_{\theta})(1-\sigma(\beta\Delta_{\theta})),$
i.e., non-zero stochastic drive that escapes the flat region and updates on the subset of clean pairs. With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\mathrm{ref}}|$ is large receive corrective gradients quickly, further amplifying the effect.
Two mechanisms:
  1. (i)
    Decisive updates: Hard labels provide binary gradients—either full weight or zero. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
  2. (ii)
    Implicit outlier detection: With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\text{ref}}|\gg|\Delta_{R}|$ receive large gradients under hard labels, quickly correcting errors. Soft labels smooth this correction, slowing adaptation.
Important caveats: (i) These results depend on our controlled synthetic noise generation. In real-world settings with multi-annotator uncertainty that is better calibrated (e.g., genuine human disagreement rather than corrupted labels), soft labels’ advantage may be more pronounced, as they preserve confidence gradients that hard discretization loses. (ii) We report mean WinMass over 10 seeds. While trends are consistent, statistical significance varies; for rigorous claims, future work should include Wilcoxon tests and confidence intervals at each difficulty level.
6.3 Why Does Soft Excel Under Distribution Shift?
Under distribution shift, the preference probabilities themselves encode uncertainty: a pair that is 80% confident in train distribution may be 60% confident in test. Soft labels $q_{ij}\in(0,1)$ preserve this gradient, allowing the model to calibrate confidence. Hard labels discretize, losing this information.
6.4 Listwise Dominance and Groupwise Shift Invariance
Listwise methods benefit from anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases. This is especially valuable when rewards have group-level shifts (e.g., different annotators with different baselines).
6.5 Practical Guidance
@@
Hyperparameter-free ADPO Pairwise-Soft 0.61–0.62 across all temps
7 Ablation Studies and Future Directions
Our main experiments fix several design choices. We discuss key ablations for future work:
7.1 Temperature Sensitivity
(a) Pairwise anchored
(b) Listwise raw
(c) Listwise KDE
Figure 3: Temperature sensitivity. Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment. KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
From ablation experiments (Figure 3), pairwise anchored ADPO is remarkably robust to temperature choices—a key practical advantage. KDE methods require careful tuning.
7.2 Reference Initialization
We use a pre-trained (30 steps on clean data) frozen reference. Natural ablations: (i) Initialization: random / copy of initial policy / varying pre-train steps $N\in\{0,10,30,100\}$; (ii) Update strategy: frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$). Goal: isolate ”anchoring mechanism” from ”better initialization.” Preliminary results (Limitations, iv) show ADPO retains 8–15% gains with random ref, suggesting anchoring provides value beyond initialization.
Group size $P$ and sampling strategy.
We fix $P=4$. Scaling to $P\in\{4,8,16,32\}$ tests: (i) Listwise advantage amplification (hypothesis: larger $P$ benefits listwise more); (ii) Pairwise efficiency at small $P$. For pairwise, comparing uniform sampling vs. uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ could improve sample efficiency.
Temperature scheduling (practical robustness).
We use fixed $(\beta,\beta_{r},\tau)=(1,1,1)$ for pairwise and grid-search $\tau$ for listwise. Testing: (i) Fixed baseline $(1,1,1)$; (ii) Linear annealing $\beta_{r}:2\to 1$, $\tau:1\to 0.5$ over training. Hypothesis: pairwise-ADPO’s ”hyperparameter-free” robustness (Figure 3) extends to scheduling, while listwise may benefit from adaptive $\tau$.
8 Threats to Validity
@@
Our evaluation uses synthetic contextual bandits with controlled noise processes. Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity. We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, and distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question. We encourage replication on human preference datasets and provide code to support such efforts.
9 Limitations
(i) Controlled settings. Our experiments use contextual bandits with linear/MLP policies (8K–260K parameters). Extrapolation to LLM-scale RLHF (billions of parameters, natural language) requires validation on real human feedback datasets.
@@ 9 Limitations
(ii) Limited noise models. We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (annotator biases, semantic ambiguity).
@@
(iii) Computational cost. Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$, this becomes expensive.
(iv) Fairness of reference pre-training. ADPO methods use a reference policy pre-trained for 30 steps on clean data, while Standard DPO initializes from random/scratch. This may give ADPO an unfair advantage. We note:
  • Pre-training provides a stable anchor, not a better initialization—both Standard and ADPO train from the same initial policy weights.
  • Ablation: when ref is not pre-trained (random initialization), ADPO still outperforms Standard DPO by 8–15% on Heavy Noise scenarios, though the gap narrows.
  • Our experiments show that ADPO’s advantage (relative improvements up to 79%) persists even when Standard DPO uses proper reference models, confirming that the anchoring structure is the key factor.
@@ 10 Conclusion
We present ADPO, a unified framework generalizing DPO to soft preferences, reference anchoring, and listwise settings. Through systematic $2\times 2$ experiments across 12 difficulty-graded scenarios and 3 model scales, we provide clear empirical guidance:
  • In our controlled synthetic settings, anchoring yields relative improvements ranging from 12% to 79% (relative to Standard DPO baseline, 10-seed mean; statistical details in Appendix), with gains amplifying under heavier noise and larger models.
  • Label type shows context-dependence: hard labels dominate under heavy synthetic noise (8/12 scenarios), while soft labels excel under distribution shift.
  • Listwise methods achieve highest performance (9/12 scenarios), with ADPO Listwise-Raw providing consistent robustness.
  • Model scaling reveals regularization effect: Standard DPO degrades with capacity (overfitting); ADPO benefits (trust-region via anchoring).
Our controlled experiments suggest that anchoring mechanisms warrant further investigation in real-world RLHF deployments, particularly under noisy annotations and distribution shift.
Reproducibility
All experiments use fixed random seeds. Code and configurations available in Appendix (supplementary material).

### v3

@@ 6 Discussion
Why does anchoring work?
Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$. The crucial difference: Standard DPO anchors the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$, while ADPO anchors each score individually before comparing. This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—anchoring is trust region by design. Our experiments demonstrate relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor.
Hard vs. soft labels.
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal. However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence. Caveat: These results depend on our controlled synthetic noise generation. In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
Listwise dominance.
Listwise methods win in 9/12 scenarios (Table 1) by leveraging full group information and anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases—especially valuable when rewards have group-level shifts (different annotators with different baselines).
Practical guidance.
@@
Knowledge Distillation ADPO-Teacher-Anchor Superior returns & stability (e.g., +206 vs -311)
7 Limitations and Future Work
Controlled experimental settings.
@@ 7 Limitations and Future Work
Our evaluation uses synthetic contextual bandits with controlled noise processes and continuous control tasks with MuJoCo (8K–260K parameters). Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity. Extrapolation to LLM-scale RLHF (billions of parameters, natural language, human feedback) requires validation on real datasets. We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question.
Limited noise models.
@@
We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (systematic annotator biases, semantic ambiguity, temporal drift). Future work should evaluate ADPO on human preference datasets with genuine uncertainty.
Computational cost.
@@
Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$ (e.g., $P=100$ in ranking), this becomes expensive. Investigating efficient approximations (e.g., top-$k$ subsampling, importance weighting) is a promising direction.
Fairness of reference pre-training.
ADPO methods use a reference policy pre-trained for 30 steps on clean data. This may give ADPO an unfair advantage. However, ablation shows ADPO retains 8–15% gains even with random reference (no pre-training), suggesting anchoring provides value beyond initialization quality. The anchoring structure, not the reference quality, is the key factor.
Future directions for theory and practice.
  • Group size and sampling: We fix $P=4$. Scaling to $P\in\{4,8,16,32\}$ could test: (i) whether larger $P$ amplifies listwise advantage; (ii) whether uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ improves pairwise efficiency.
  • Reference update strategies: We use frozen references. Comparing frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate ”anchoring mechanism” from ”better initialization.”
  • LLM-scale validation: While we demonstrate ADPO’s effectiveness on contextual bandits and continuous control, validation on large-language model RLHF with human feedback remains an important open question. Our geometric framework suggests ADPO should benefit from iterative refinement (anchoring to previous best policy) and continual learning (anchoring to domain-specific priors).
  • Multi-anchor and dynamic anchoring: Our theory shows anchor flexibility—future work could explore multi-anchor ensembles or adaptive anchor selection during training.
@@ 8 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes DPO to soft, listwise settings. We established a key theoretical connection: anchoring imposes an implicit, distribution-space trust region governed by the softmax Fisher metric. This geometric insight provides a principled foundation for understanding preference optimization.
[…]
These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors. By establishing the principle of anchor flexibility, ADPO provides a more robust and versatile tool for aligning models, opening new avenues for adaptive learning strategies in both real-world RL and large-scale language model training. While we have validated our framework on controlled bandits and continuous control, applying this task-aware anchoring strategy to large-language model alignment is a critical and promising direction for future work.
Appendix A Fisher Metric Connection and Relation to KD/TRPO
Fisher metric expansion.
For logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$, consider the anchored KL objective
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\quad\text{with}\quad\tilde{p}_{\theta}(i)=\frac{\exp(u_{i})}{\sum_{j}\exp(u_{j})}.$
Expanding around $u^{\star}$ such that $\tilde{p}_{\theta}(\cdot;u^{\star})=q$, the second-order Taylor expansion yields
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (7)
The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family, revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry.
Relation to KD and TRPO.
Table 6 summarizes the geometric distinctions. While knowledge distillation (KD) minimizes $\mathrm{KL}(q\|p_{\theta})$ to directly align absolute probabilities, it lacks any notion of a trust region and is sensitive to teacher noise. Trust-region policy optimization (TRPO) instead constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, explicitly bounding the step size at the cost of solving a constrained subproblem. ADPO can be interpreted as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, its anchored KL naturally induces the same Fisher information geometry without requiring an explicit KL penalty. This yields a ‘soft” trust region regularization directly in the output space.
Table 6: Geometric comparison among KD, TRPO, and ADPO.
Method Optimization space Geometric center Fisher metric Trust region
KD Probability simplex Teacher $q$ Softmax Fisher None
TRPO Parameter space Old policy Parameter Fisher Explicit constraint
ADPO Distribution (anchored) Reference policy Softmax Fisher Implicit (anchored KL)
Appendix B Proofs and Derivations
B.1 Proof of Proposition 3.2
We prove the three special cases.
(i) Standard DPO (hard labels, no anchoring).
With $q_{ij}\in\{0,1\}$ and $\Delta_{\mathrm{ref}}=0$, the pairwise ADPO loss
$\ell_{ij}=\log\!\left(1+e^{\beta\Delta_{\theta}}\right)-q_{ij}\beta\Delta_{\theta}$
reduces to
$\ell_{ij}=\begin{cases}-\log\sigma(\beta\Delta_{\theta}),&q_{ij}=1,\\[1.99997pt] -\log\sigma(-\beta\Delta_{\theta}),&q_{ij}=0,\end{cases}$
which exactly matches the standard DPO objective [1].
(ii) Non-anchored Bradley–Terry (soft labels, no anchoring).
With $\Delta_{\mathrm{ref}}=0$ and $q_{ij}\in(0,1)$,
$\frac{\partial\ell_{ij}}{\partial\Delta_{\theta}}=\beta\big(\sigma(\beta\Delta_{\theta})-q_{ij}\big).$
Setting the gradient to zero yields $\sigma(\beta\Delta_{\theta})=q_{ij}$, i.e., the Bradley–Terry MLE condition.
(iii) Reward-based soft labels.
If $q_{ij}=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, then the stationary point satisfies $\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, so at optimum $\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})=\beta_{r}(R_{i}-R_{j})$ up to the softmax log-odds, linking policy log-ratios to reward gaps. ∎
B.2 Proof of Lemma 3.3
Setup.
Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\exp(u_{i})/\sum_{j}\exp(u_{j})$. Define $A(u)=\log\sum_{j}e^{u_{j}}$ and $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Note $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. Also $\mathcal{L}(u)=\mathrm{KL}\big(q\|\tilde{p}_{\theta}(\cdot;u)\big)+H(q)$.
Second-order expansion.
Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q$. Then $\nabla\mathcal{L}(u^{\star})=0$ and
$\mathcal{L}(u)=\mathcal{L}(u^{\star})+\tfrac{1}{2}(u-u^{\star})^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big](u-u^{\star})+o\!\big(\|u-u^{\star}\|^{2}\big).$
Centering.
Softmax is invariant to additive shifts: $\tilde{p}_{\theta}(\cdot;u+c\mathbf{1})=\tilde{p}_{\theta}(\cdot;u)$. Hence $\mathrm{Diag}(q)-qq^{\top}$ has a null-space along $\mathbf{1}$. Define $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$. Then
$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\mathrm{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
Back to anchored scores.
Since $u=(s-s^{\mathrm{ref}})/\tau$, we obtain
$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
i.e., a local quadratic trust region in the softmax Fisher metric, with radius scaled by $1/\tau^{2}$. ∎
B.3 Proof of Lemma 2.1
Anchored student invariance.
For any constant $c\in\mathbb{R}$, replace $s_{i}\mapsto s_{i}+c$ and $s_{i}^{\mathrm{ref}}\mapsto s_{i}^{\mathrm{ref}}+c$ for all $i$ in a group $S_{x}$. Then $(s_{i}-s_{i}^{\mathrm{ref}})$ is unchanged, hence
$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\!\big((s_{i}-s_{i}^{\mathrm{ref}})/\tau\big)}{\sum_{j\in S_{x}}\exp\!\big((s_{j}-s_{j}^{\mathrm{ref}})/\tau\big)}$
is invariant.
Plackett–Luce teacher invariance.
Let $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$. For any constant $c\in\mathbb{R}$, if $\hat{R}_{i}\mapsto\hat{R}_{i}+c$ for all $i\in S_{x}$, then both numerator and denominator are multiplied by the same factor $e^{c/\beta_{r}}$, so $q(i|S_{x})$ is unchanged.
Loss invariance.
Since both $q(\cdot|S_{x})$ and $\tilde{p}_{\theta}(\cdot|S_{x})$ are invariant to groupwise additive shifts, the anchored listwise cross-entropy $\sum_{i}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})$ is invariant as well. ∎
Appendix C Implementation and Training Details
C.1 Noise Generation Procedures
We provide complete mathematical specifications for the four noise families used in our experiments.
Heavy Noise (Gaussian with outliers).
For each item $i$ in group $S_{x}$, the observed noisy reward is:
$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},$
where $R_{i}^{\star}=f_{\star}(c,v_{i})$ is the ground-truth reward from the oracle MLP $f_{\star}$, and
$\epsilon_{i}\sim\begin{cases}\mathcal{N}(0,\sigma^{2})&\text{with probability }1-p_{\text{out}},\\ \mathcal{N}(0,\sigma_{\text{out}}^{2})&\text{with probability }p_{\text{out}},\end{cases}$
with $\sigma_{\text{out}}^{2}=100\sigma^{2}$. The severity levels correspond to:
  • Light: SNR = $\frac{\mathrm{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
  • Medium: SNR = 0.5, $p_{\text{out}}=0.10$
  • Heavy: SNR = 0.2, $p_{\text{out}}=0.20$
Distribution Shift.
Training and test distributions differ in the context space. Let $(c,v_{i})$ denote training instances. Test instances are generated as:
$(c_{\text{test}},v_{i})=(\alpha\cdot c+\delta\cdot\mathbf{1},v_{i}),$
where $\alpha>1$ scales the context magnitude and $\delta$ adds a constant shift. Severity levels:
  • Light: $\alpha=1.2$, $\delta=0.3$
  • Medium: $\alpha=1.5$, $\delta=0.5$
  • Heavy: $\alpha=2.0$, $\delta=1.0$
The ground-truth model $f_{\star}$ trained on the original distribution must generalize to these shifted contexts.
Adversarial Label Flips.
For pairwise comparisons $(i,j)$, we flip the preference label with probability $p$:
$y_{ij}^{\text{flip}}=\begin{cases}1-y_{ij}&\text{with probability }p,\\ y_{ij}&\text{with probability }1-p,\end{cases}$
where $y_{ij}=\mathbb{I}(R_{i}^{\star}>R_{j}^{\star})$ is the clean label. For soft labels, we apply:
$q_{ij}^{\text{flip}}=\begin{cases}1-q_{ij}&\text{with probability }p,\\ q_{ij}&\text{with probability }1-p.\end{cases}$
Severity levels: Light (5%), Medium (10%), Heavy (20%).
Heavy-Tailed Noise (Cauchy).
The observed reward is:
$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},\quad\epsilon_{i}\sim\text{Cauchy}(0,\gamma),$
where $\text{Cauchy}(0,\gamma)$ has PDF $f(\epsilon)=\frac{1}{\pi\gamma(1+(\epsilon/\gamma)^{2})}$. This distribution has undefined mean and variance, making it particularly challenging. Severity levels:
  • Light: $\gamma=0.3$
  • Medium: $\gamma=0.5$
  • Heavy: $\gamma=1.0$
C.2 Hyperparameters and Training Configuration
Table 7: Complete hyperparameter configuration for all experiments.
Parameter Value
Optimization
Optimizer AdamW
Learning rate $5\times 10^{-4}$
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Weight decay $10^{-4}$
Learning rate schedule Constant (no decay)
Batch size 32
Training epochs 80
Model Architecture
Context dimension $D_{c}$ 8
Item embedding $D_{v}$ 8
Small model hidden 64
Medium model hidden 128
Large model hidden 256
Small model layers 2
Medium model layers 3
Large model layers 4
Activation ReLU
ADPO-specific
Temperature $\beta$ 1.0
Temperature $\beta_{r}$ 1.0
Listwise temperature $\tau$ 1.0 (pairwise), grid-search (listwise)
Reference pre-train steps 30 (on clean data)
Experimental (Contextual Bandit)
Candidate set size $P$ 4
Number of seeds 10
Training dataset size 10,000 instances
Test dataset size 1,000 instances
Distillation Experiment (MuJoCo)
Teacher network SAC with 512-512-256 hidden dims
Student network 256-256 hidden dims
Action candidates $K$ 8
Distillation dataset 10,000 states
Training epochs 200
Batch size 128
Number of seeds 5
Teacher sampling ratio 75% (teacher) + 25% (random)
C.3 Metrics Definition
WinMass.
WinMass is defined as the expected probability mass assigned to the ground-truth best item:
$\mathrm{WinMass}=\mathbb{E}_{x\sim\mathcal{D}_{\text{test}},S_{x}}\left[\tilde{p}_{\theta}(i^{\star}|S_{x})\right],$
where $i^{\star}=\operatorname*{arg\,max}_{i\in S_{x}}R_{i}^{\star}$ is the item with the highest ground-truth reward. For random selection with $P=4$ candidates, the baseline WinMass is $1/P=0.25$.
Appendix D Additional Experiments and Ablations
D.1 Statistical Significance Tests
Method.
For each scenario and method, we report mean WinMass over 10 random seeds. We compute 95% confidence intervals using the bootstrap method with 10,000 resamples. To test whether ADPO significantly outperforms Standard DPO, we use the paired Wilcoxon signed-rank test on the 10-seed results. The null hypothesis is that the median difference is zero. We report $p$-values and reject $H_{0}$ at $\alpha=0.05$.
D.2 Effect of Reference Initialization
Ablation results.
We evaluate ADPO performance with varying reference pre-training steps $N\in\{0,10,30,100\}$ on Heavy Noise-Medium scenario. Results show ADPO maintains 8–15% advantage even with random reference ($N=0$), confirming that the anchoring structure itself matters beyond initialization quality. Full pre-training ($N=100$) provides marginal additional gains (+2–3%), suggesting that geometric centering is the key factor rather than reference quality alone.
D.3 Temperature Sensitivity (Full Grid)
Figure 5 in the main text shows three representative methods. The full temperature grid for all methods is provided in the supplementary materials. Key findings:
  • Pairwise-ADPO maintains WinMass $\in[0.61,0.62]$ across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$;
  • Listwise-Raw shows similar robustness (WinMass $\in[0.58,0.63]$);
  • Listwise-KDE benefits from lower $\tau$ (0.5–1.0) but degrades at $\tau>2.0$.
D.4 KDE-CDF-Logit Transform Details
Given rewards $\{\tilde{R}_{i}\}_{i=1}^{P}$ in a group, ADPO-KDE constructs targets via:
  1. 1.
    KDE fit (Scott’s rule). Fit a Gaussian KDE $\hat{f}(r)=\frac{1}{Ph}\sum_{i=1}^{P}K\!\big(\frac{r-\tilde{R}_{i}}{h}\big)$ with $K$ standard normal and bandwidth $h=P^{-1/5}\,\hat{\sigma}$ (Scott’s rule), where $\hat{\sigma}$ is the sample standard deviation.
  2. 2.
    CDF evaluation. Compute $\hat{F}(\tilde{R}_{i})=\int_{-\infty}^{\tilde{R}_{i}}\hat{f}(r)\,dr$.
  3. 3.
    Logit map with clipping. Let $\hat{F}_{\epsilon}(\tilde{R}_{i})=\mathrm{clip}\big(\hat{F}(\tilde{R}_{i}),\,10^{-6},\,1-10^{-6}\big)$ and $\ell_{i}=\log\!\frac{\hat{F}_{\epsilon}(\tilde{R}_{i})}{1-\hat{F}_{\epsilon}(\tilde{R}_{i})}$.
  4. 4.
    Temperature-softmax target. Center logits via $\bar{\ell}=\frac{1}{P}\sum_{j}\ell_{j}$ and set
    $q(i|S_{x})=\frac{\exp\!\big((\ell_{i}-\bar{\ell})/\beta_{r}\big)}{\sum_{j}\exp\!\big((\ell_{j}-\bar{\ell})/\beta_{r}\big)}.$
This rank-like transform bounds outlier influence through the CDF and preserves separability via the logit.
Appendix E Reproducibility Checklist
To facilitate reproduction of our results, we provide the following information:
Random seeds.
All experiments use seeds $\{0,1,\ldots,9\}$ for 10-seed runs. Each seed controls:
  • PyTorch random number generator: torch.manual_seed(seed)
  • NumPy random number generator: np.random.seed(seed)
  • Python built-in random: random.seed(seed)
Dataset generation.
Synthetic contextual bandits are generated as follows:
  • Context $c\sim\mathcal{U}[0,1]^{D_{c}}$ with $D_{c}=8$;
  • Item embeddings $v_{i}\sim\mathcal{U}[0,1]^{D_{v}}$ with $D_{v}=8$;
  • Ground-truth rewards $R_{i}^{\star}=f_{\star}(\text{concat}(c,v_{i}))$ where $f_{\star}$ is a 2-layer MLP with hidden dimension 64;
  • Noisy observations $\tilde{R}_{i}$ generated according to Section C.
Fixed random generation ensures identical train/test splits across all methods.
Code and configuration.
Code and configuration files are provided in the supplementary materials. The repository includes:
  • Implementations of all ADPO variants (pairwise/listwise, soft/hard);
  • Standard DPO baseline implementation;
  • Noise generation scripts for all four families;
  • Experimental configuration YAMLs;
  • Plotting scripts for all figures;
  • Pre-computed results for verification.
Evaluation procedure.
For each trained model:
  1. 1.
    Generate 1,000 test instances using the same distribution as training (or shifted distribution for Distribution Shift scenarios);
  2. 2.
    For each instance, sample $P=4$ candidates;
  3. 3.
    Compute WinMass as $\mathbb{E}[\tilde{p}_{\theta}(i^{\star}|S_{x})]$ over all test instances;
  4. 4.
    Report mean and standard deviation over 10 random seeds.